# Deep Linear Hawkes Processes

### v1

@@ C.3 Dataset Pre-processing
We use the default train/validation/test splits for EasyTPP benchmark datasets. For MIMIC-II, we copy Du et al. (2016) and keep the 325 test sequences in the test split, and further split the 2,935 training sequences into 2,600 for training and 325 for validation. In our pre-processed datasets, Last.fm and EHRSHOT, we randomly partition into subsets containing 70%, 15%, 15% of all sequences for training/validation/test respectively. We provide a high-level description of all the datasets we used, followed by our pre-processing procedure of Last.fm and EHRSHOT in more detail. Note that for datasets that contain concurrent events or effectively discrete times, we apply a small amount of jittering to ensure no modeling assumptions are violated in the MTPP framework.
@@
Amazon (Ni et al., 2019) contains user product reviews where product categories are considered as marks. Retweet (Zhao et al., 2015) predicts the popularity of a retweet cascade, where the event type is decided by if the retweet comes from users with “small”, “medium”, or “large” influences, measured by number of followers (Mei & Eisner, 2017). Taxi data (Whong, 2014; Mei et al., 2019) uses data from the pickups and dropoffs of New York taxi and the marks are defined as the Cartesian product of five discrete locations and two actions (pickup/dropoff). Taobao (Xue et al., 2022) describes the viewing patterns of users on an e-commerce site, where item categories are considered as marks. StackOverflow contains the badges (defined as marks) awarded to users on a question-answering website. Finally, MIMIC-II (Saeed et al., 2002) records different diseases (used as marks) during hospital visits of patients. We add a small amount of noise to the MIMIC-II event times so that events do not lie on a fixed grid. Both StackOverflow and MIMIC-II datasets were first pre-processed by Du et al. (2016).
@@
Last.fm Celma Herrada et al. (2009); McFee et al. (2012) records 992 users’ music listening habits that has been widely used in MTPP literature (Kumar et al., 2019; Boyd et al., 2020; Bosser & Taieb, 2023). Mark types are defined as the genres of a song, and each event is a play of a particular genre. Each sequence represents the monthly listening behavior of each user, with sequence lengths between 5 and 500. If the song is associated with multiple genres we select a random one of the genres, resulting in a total of 120 different marks.
@@
EHRSHOT Wornow et al. (2023) is a newly proposed large dataset of longitudinal de-identified patient medical records, and has rich information such as hospital visits, procedures, and measurements. We introduce an MTPP dataset derived from EHRSHOT, where medical services and procedures are treated as marks, as identified by Current Procedural Terminology (CPT-4) codes. Each patient defines an event sequence, and we retain only CPT-4 codes with at least 100 occurrences in the dataset. For the $<1\%$ events of events where there are more than 10 codes at a single timestamp, we retain the top 10 codes with the most frequencies and discard the rest. We then add a small amount of random noise to the event time to ensure they are not overlapping. This process ensures we still satisfy the MTPP framework, and can reasonably instead compute top-10 accuracy for the next mark prediction. Other work has considered extending the MTPP framework to consider simultaneous event occurrence (Chang et al., 2024). Then we standardize each sequence to start and end with start and end of a sequence events. Note that we do not score these events. Event times are normalized to be in hours. We discard sequences that have less than 5 events and a single timestamp. This leads to the final version of our dataset to have 668 marks, and the sequence lengths range from 5 to 3955 events, reflecting patient histories that can span multiple years.
@@ D.1 Full Results on Benchmark Datasets
We provide the full log-likelihood results and corresponding plots in Table 7 and Fig. 5 respectively, where we decompose the likelihood into time and mark likelihoods. The improvement of our DLHP model is mainly driven by better modeling of time, though we also often obtain best- or second-best predictive performance on marks from the next event prediction accuracy results conditioned on true event time in Table 8. In aggregate, our model achieves a 1.38 per-event likelihood ratio between itself and the next best method across all datasets (a 38% improvement in likelihood). This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating. Doing so is equivalent to taking the geometric mean across likelihood ratios.
We also include root mean square error (RMSE) in Table 9 as one of the commonly used metrics in MTPP literature, with the caveat that it purely describes how well the model captures the expected value of the next-event time. In other words, it is not informative when the inter-event times have a multimodal distribution – which we often encounter in practice with real data. Intensity-based methods were sampled using the thinning algorithm and averaged over 64 sampled candidate times when memory allows. For our method, we include a simple linear probe on the output hidden vector to directly estimate the next event time, which is substantially faster in wallclock runtime than sampling even when accounting for fine-tuning the linear probe parameters. For RMSE, IFTPP is the best performer overall; however, across intensity-based methods we find our model to be competitive. We posit this is due to the simplified parameterization of IFTPP lending itself well to capturing the expected time of the next event.
Table 7: Complete per event log-likelihood results on the held-out test for the eight benchmark datasets we consider, averaged over 5 random seeds. In Table 7(a) we show the full log-likelihood. We then decompose this log-likelihood into the log-likelihood of the event time in Table 7(b), and the time-conditional log-likelihood of the mark type in Table 7(c). OOM indicates out of memory; standard deviation in parentheses. We highlight the best-performing model in bold and underline the second-best. We also report the average rank of models across datasets as a summary metric (lower is better). DLHP is consistently the best or second best-performing model across all datasets.
(a) Full log-likelihood results (equal to the summation of Table 7(b) and Table 7(c)). Extended version of Table 1.
@@
IFTPP -1.988 (0.001) -0.844 (0.007) -0.282 (0.001) -1.391 (0.005) -1.571 (0.003) -1.769 (0.004) -0.239 (0.002) -3.956 (0.192) 3.8
DLHP (Ours) -1.871 (0.002) -0.767 (0.000) -0.211 (0.002) -1.415 (0.005) -1.521 (0.008) -1.814 (0.025) -0.145 (0.040) -2.893 (0.009) 1.9
Figure 5: Visualization of $\mathcal{L}_{\text{Total}}$ decomposed into $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$ for all models and all datasets relative to RMTPP, normalized by number of events, as discussed in Section 5.2. The improvement of DLHP is mainly driven by better modeling of $\mathcal{L}_{\text{Time}}$.
Table 8: Next event prediction accuracy (reported as a percentage, higher is better) conditioned on the true event time. We report top 1 accuracy for all datasets except for top 10 accuracy for EHRSHOT, due to the pre-processing procedure described in Section C.3. We bold the best result per dataset, and underline the runner-up.
Model Next Mark Accuracy (%) ($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT (Top 10)
RMTPP 30.89 53.76 91.32 60.78 46.03 52.42 92.32 36.79 6.1
NHP 39.32 61.43 92.97 61.48 47.21 56.55 94.32 76.41 1.8
SAHP 31.45 54.97 92.23 60.45 46.54 51.82 87.42 52.24 5.5
THP 34.63 60.20 91.46 59.99 46.79 53.38 91.09 51.85 5.3
AttNHP 38.88 60.90 92.74 61.32 48.38 56.39 93.65 OOM 2.6
IFTPP 35.90 54.92 91.76 60.93 45.66 56.44 93.32 66.38 4.4
DLHP (Ours) 40.76 61.36 93.27 61.19 47.86 56.28 96.55 79.20 1.9
Table 9: Comparison of RMSE of the next-event time prediction on the benchmarks we consider in Section 5.2 in the same format as Table 8. Lower RMSE values indicate better performance.
Model Next event time RMSE ($\downarrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP 0.361 16152 0.284 0.127 1.054 15.864 0.754 3445 4.8
NHP 0.342 15322 0.283 0.127 1.028 15.841 0.739 3430 3.1
SAHP 0.345 16018 0.285 0.126 1.034 15.830 0.805 3418 3.8
THP 0.335 15848 0.286 0.126 1.029 15.878 0.781 3504 4.3
AttNHP 1.214 16220 1.273 0.130 1.311 15.889 0.852 OOM 7.0
IFTPP 0.332 2568 0.280 0.126 0.975 15.550 0.713 1899 1.0
DLHP (Ours, sampling) 0.395 15223 0.282 0.127 1.021 15.844 0.764 3421 3.4
DLHP (Ours, linear probe) 0.337 14642 0.283 0.128 1.038 15.710 0.810 3350 N/A
@@ D.2 Full Results for Synthetic Poisson Experiments
We present the full results in Fig. 6 for all models regarding the synthetic experiments discussed in Section 5.1. All models are trained until convergence using a set of 5,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training to accommodate the sparsity of events. We used small models so they do not overfit; model architecture and parameter counts are reported in Table 10. We plot the background intensity conditioned on empty sequences using 1,000 equidistant grid points between the start and end points. Our model is the only one that perfectly recovers the underlying ground truth intensity, while also using the fewest parameters.
Figure 6: Results for all baseline models for the synthetic Poisson experiment introduced in Section 5.1. The estimated intensity (blue lines) conditioned on an empty sequence are plotted against the ground truth (dotted black lines).
@@ D.3 Additional Synthetic Results on Multivariate Hawkes Processes
We evaluate our model and baseline models against the true model on a randomly initiated parametric Hawkes process with three possible marks. Following the notation in Section 2.1, we draw all parameters from the following distributions: $\bm{\nu}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.1,0.5]$, $\bm{\alpha}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.5,0.8]$, and $\bm{\beta}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.4,1.2]$ for $i,j\in\{1,2,3\}$.
All models are trained until convergence using a set of 50,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training. Model architecture and parameter counts are reported in Table 11. We plot three example sequences drawn for an additional test set for each model in Figs. 7 and 8, using 1,000 equidistant grid points for any inter-event interval. Dotted lines refer to the intensities under the true underlying parametric model; solid lines are different model estimates from trained models.
As we see in inhomogeneous Poisson processes, our model can recover the ground truth intensities with the fewest parameters. Both neural Hawkes processes and our DLHP show almost perfect recovery of parametric Hawkes processes, especially before seeing any event happening, and at event times. It is also worth noting that our model is 7-9$\times$ quicker than NHP and AttNHP regarding wall-clock runtime on a single A5000 GPU. Our results on synthetic experiments validate the model’s ability to recover the ground truth intensities.
We further evaluate all models quantitatively using 1,000 test sequences generated from the same multivariate Hawkes process and evaluated both log-likelihood and RMSE for the immediate next event. We see our method competitive again on both metrics.
Table 11: Model architectures and parameter counts for parametric Hawkes processes experiments.
@@
IFTPP $h=16$ 1965
DLHP (Ours) $h=8,p=4,l=2$ 358
Figure 7: Our proposed DLHP model trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
Figure 8: Baseline models trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
Table 12: Performance comparison of models on the synthetic Hawkes process experiment presented above. Higher per event log-likelihood indicates better performance, whereas lower root mean squared error (RMSE) indicates better performance.
@@
RMTPP -0.550 0.570
NHP -0.530 0.565
SAHP -0.537 0.569
THP -0.543 0.570
AttNHP -0.533 0.655
IFTPP -0.534 0.540
DLHP (Ours) -0.527 0.566
D.4 Ablation for Different DLHP Variants
@@
We perform an ablation study of different model variants that we proposed on all datasets and summarize the results in Table 13. We train EHRSHOT using 10% of its training data because larger dataset scale requires more training time (but use the original validation and test sets for model selection and reporting results). Forward and backward discretization are very close in performance, with backwards discretization having a slight edge. Models that are input-dependent achieve better performance on most datasets, although on certain datasets input dependence appears to harm performance. It is an interesting direction for future work to explore theoretically and empirically when each of these variants is best. We select backward discretization with input dependence for the results in the main paper.
@@
Table 13: Ablation for different model variants log-likelihood (LL). ID stands for input-dependent, see Section 3.4. Backward and Forward respectively refer to using $\mathbf{u}_{t_{i-1}}$ and $\mathbf{u}_{t_{i}-}$ (i.e., the previous right limit or current left limit), see Section B.4.
Dataset Model variant LL Arrival time LL Mark LL conditioned on time
@@
D.5 Model Calibration
To further probe the models, we evaluate the calibration metrics of MTPPs that are proposed in literature (Bosser & Taieb, 2023), which has a different focus than log-likelihood-based evaluation. On a high level, calibration describes how well the uncertainty in the model is reflected in the observed data. However, a model can achieve perfect calibration by predicting the marginal distribution, so better calibration does not necessarily transform into better predictive performance. We therefore present these metrics as a secondary metric (secondary to log-likelihood per Daley & Vere-Jones (2003)) for investigating the performance of different models. We provide summarized statistics for both probabilistic calibration error (PCE) for time calibration and expected calibration error (ECE) for mark calibration in Table 14, and visualize the calibration curves in Figs. 9 and 10. From our results, all MTPP models are well-calibrated on most of the datasets, especially on mark predictions.

### v2

@@ C.3 Dataset Pre-processing
We use the default train/validation/test splits for EasyTPP benchmark datasets. For MIMIC-II, we copy Du et al. [2016] and keep the 325 test sequences in the test split, and further split the 2,935 training sequences into 2,600 for training and 325 for validation. In our pre-processed datasets, Last.fm and EHRSHOT, we randomly partition into subsets containing 70%, 15%, 15% of all sequences for training/validation/test respectively. We provide a high-level description of all the datasets we used, followed by our pre-processing procedure of Last.fm and EHRSHOT in more detail. Note that for datasets that contain concurrent events or effectively discrete times (e.g., StackOverflow, Retweet), we apply a small amount of jittering to ensure no modeling assumptions are violated in the MTPP framework.
@@
Amazon [Ni et al., 2019] contains user product reviews where product categories are considered as marks. Retweet [Zhao et al., 2015] predicts the popularity of a retweet cascade, where the event type is decided by if the retweet comes from users with “small”, “medium”, or “large” influences, measured by number of followers [Mei and Eisner, 2017]. Taxi data [Whong, 2014, Mei et al., 2019] uses data from the pickups and dropoffs of New York taxi and the marks are defined as the Cartesian product of five discrete locations and two actions (pickup/dropoff). Taobao [Xue et al., 2022] describes the viewing patterns of users on an e-commerce site, where item categories are considered as marks. StackOverflow contains the badges (defined as marks) awarded to users on a question-answering website. Finally, MIMIC-II [Saeed et al., 2002] records different diseases (used as marks) during hospital visits of patients. We add a small amount of noise to the MIMIC-II event times so that events do not lie on a fixed grid. Both StackOverflow and MIMIC-II datasets were first pre-processed by Du et al. [2016].
@@
Last.fm Celma Herrada et al. [2009], McFee et al. [2012] records 992 users’ music listening habits that has been widely used in MTPP literature [Kumar et al., 2019, Boyd et al., 2020, Bosser and Taieb, 2023]. Mark types are defined as the genres of a song, and each event is a play of a particular genre. Each sequence represents the monthly listening behavior of each user, with sequence lengths between 5 and 500. If the song is associated with multiple genres we select a random one of the genres, resulting in a total of 120 different marks.
@@
EHRSHOT Wornow et al. [2023] is a newly proposed large dataset of longitudinal de-identified patient medical records, and has rich information such as hospital visits, procedures, and measurements. We introduce an MTPP dataset derived from EHRSHOT, where medical services and procedures are treated as marks, as identified by Current Procedural Terminology (CPT-4) codes. Each patient defines an event sequence, and we retain only CPT-4 codes with at least 100 occurrences in the dataset. For the $1\%$ events of events where there are more than 10 codes at a single timestamp, we retain the top 10 codes with the most frequencies and discard the rest. We then add a small amount of random noise to the event time to ensure they are not overlapping. This process ensures we still satisfy the MTPP framework, and can reasonably instead compute top-10 accuracy for the next mark prediction; other work has considered extending the MTPP framework to consider simultaneous event occurrence [Chang et al., 2024]. Then we standardize each sequence to start at $t=0$ and pad the start and end of sequences with a specific padded event token. Note that we do not score these events. Event times are normalized to be in hours. We discard sequences that have less than 5 events or a single timestamp. This leads to the final version of our dataset having 668 marks, and the sequence lengths range from 5 to 3955 events, reflecting patient histories that can span multiple years.
We include code in our forked repository for preparing the EHRSHOT event sequence dataset from the raw EHRSHOT dataset. Note that we cannot distribute the raw data (or derivative dataset) under the terms of the original EHRSHOT dataset requiring credentialed access through PhysioNet.
@@ D.1 Full Results on Benchmark Datasets
We provide the full log-likelihood results and corresponding plots in Table˜8 and Fig.˜6 respectively, where we decompose the likelihood into time and mark likelihoods. The improvement of our S2P2 model is mainly driven by better modeling of time, though we also often obtain best- or second-best predictive performance on marks from the next event prediction accuracy results conditioned on true event time in Table˜2(c). In aggregate, our model achieves a 1.33 per-event likelihood ratio between itself and the next best method across all datasets (a 33% improvement in likelihood). This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating. Doing so is equivalent to taking the geometric mean across likelihood ratios.
Configuration and training details of all models can be found in Section˜C.1. As discussed in Section˜7 and grouped in the results, models with continuous-time hidden states can present a richer class of intensities and often empirically outperform those with discrete hidden states. Note for the RMSE results in Table˜2(b), we follow Mei and Eisner [2017] and use the expected next event time as next event time predictions to minimize the Bayes risk. Unlike them, however, we estimate these with the trapezoidal rule rather than Monte-Carlo simulation via the thinning algorithm. In practice, we have found this to produce an estimator with much lower variance and be faster due to being more readily parallelizable. This stands in contrast to the thinning algorithm which has more hyperparameters (e.g., dominating rate, sampling boundary) that can exacerbate bias.
Table 8: Complete per event log-likelihood results on the held-out test for the eight benchmark datasets we consider, averaged over 5 random seeds. In Table˜8(a) we show the full log-likelihood. We then decompose this log-likelihood into the log-likelihood of the event time in Table˜8(b), and the time-conditional log-likelihood of the mark type in Table˜8(c). OOM indicates out of memory; standard deviation in parentheses. We box the best-performing model and underline the second-best. We also report the average rank of models across datasets as a summary metric (lower is better). S2P2 is consistently the best or second best-performing model across all datasets.
(a) Full log-likelihood results (equal to the summation of Table˜8(b) and Table˜8(c)). Extended version of Table 2(a).
@@
S2P2 (Ours) -1.871 (0.002) -0.767 (0.000) -0.211 (0.002) -1.415 (0.005) -1.521 (0.008) -1.814 (0.025) -0.131 (0.014) -2.893 (0.009) 1.9
Figure 6: Visualization of $\mathcal{L}_{\text{Total}}$ decomposed into $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$ for all models and all datasets relative to RMTPP, normalized by number of events, as discussed in Section˜6. The improvement of S2P2 is mainly driven by better modeling of $\mathcal{L}_{\text{Time}}$, while it improves both $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$.
@@ D.2 Full Results for Synthetic Poisson Experiments
We present the full results in Fig.˜7 for all models regarding the synthetic Poisson experiments discussed in Section˜5. All models are trained until convergence using a set of 5,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training to accommodate the sparsity of events. We used small models so they do not overfit; model architecture and parameter counts are reported in Table˜9. We plot the estimated intensity conditioned on empty sequences using 1,000 equidistant grid points between the start and end points. Our model is the only one that perfectly recovers the underlying ground truth intensity, while also using the fewest parameters.
Figure 7: Results for all baseline models for the synthetic Poisson experiment introduced in Section 5. The estimated intensity (blue lines) conditioned on an empty sequence are plotted against the ground truth (dotted black lines).
@@ D.3 Additional Synthetic Results on Multivariate Hawkes Processes
We evaluate our model and baseline models against the true model on a randomly initiated parametric Hawkes process with three possible marks. Following the notation in Section˜2.1, we draw all parameters from the following distributions: $\bm{\nu}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.1,0.5]$, $\bm{\alpha}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.5,0.8]$, and $\bm{\beta}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.4,1.2]$ for $i,j\in\{1,2,3\}$.
All models are trained until convergence using a set of 50,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training. Model architecture and parameter counts are reported in Table˜10. We plot three example sequences drawn for an additional test set for each model in Fig.˜8, using 1,000 equidistant grid points for any inter-event interval. Dotted lines refer to the intensities under the true underlying parametric model; solid lines are different model estimates from trained models.
[…]
Table 10: Model architectures and parameter counts for multivariate Hawkes processes experiments.
@@
S2P2 (Ours) $h=8,p=4,l=2$ 358
Table 11: Performance comparison of models on the multivariate Hawkes processes experiment presented above. Higher per-event log-likelihood indicates better performance, whereas lower root mean squared error (RMSE) indicates better performance.
@@
RMTPP -0.550 0.648
SAHP -0.537 0.647
THP -0.543 0.648
IFTPP -0.534 0.647
MHP -0.551 0.648
NHP -0.530 0.647
AttNHP -0.533 0.652
S2P2 (Ours) -0.527 0.647
Figure 8: Our proposed S2P2 model and baseline models trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process where $K=3$. For each model, three example test sequences are plotted; the locations of colored bars $|||$ indicate the true event times.
D.4 Full Results for Hawkes and Self-Correcting Process
We train our approach on synthetic data generated from known, classical temporal point processes, namely the self-correcting and Hawkes processes. These are characterized by intensity functions $\lambda_{t}^{\text{SC}}=\exp\left(t-0.5N_{t}\right)$ and $\lambda_{t}^{\text{H}}=0.5+\sum_{i=1}^{N_{t}}0.5\exp(t_{i}-t)$, respectively. Models were fit using data drawn from these processes, 6,000 sequences for training and 2,000 for validation. The learned intensity functions, evaluated on a held-out test sequence, can be seen for the self-correcting process data in Fig.˜9 and for the Hawkes process data in Fig.˜10 for all methods. The hyperparameters for the models were chosen by a grid search (see Table˜12). We see many of the models, including ours, do well at capturing the ground truth intensity.
Figure 9: Synthetic self-correcting process experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence. The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
Figure 10: Synthetic Hawkes process experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence. The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
Table 12: Model architectures and parameter counts for synthetic Hawkes and self-correcting process experiments.
Model Architecture # Parameters
RMTPP $h=16$ 627
SAHP $h=16,l=3,\mathrm{heads}=4$ 2554
THP $h=16,l=3,\mathrm{heads}=4$ 2500
IFTPP $h=32$ 6859
MHP $h=16$ 13556
NHP $h=32$ 14658
AttNHP $h=16,t=16,l=2,\mathrm{heads}=4$ 13298
S2P2 (Ours) $h=16,p=16,l=4$ 7202
D.5 Full Results for Long-Range Dependency Experiment
To measure the ability to capture long-range dependencies by neural MTPPs, we constructed a generative process with long-range dependencies. For this, we generate sequences over the time window of $[0,100]$, with three possible marks. The first mark is a “distractor” mark, meaning it has no influence over other events. These events are drawn from a homogeneous Poisson process with rate 1. The second mark is a “trigger” mark, which are directly tied to the third “target” mark. Triggers are also drawn from a homogeneous Poisson process with rate 0.1. For every trigger event $(t_{i},k_{i}=2)$ drawn, a corresponding target event $(t_{j},k_{j}=3)$ is generated conditionally independent of all other events according to $t_{j}|t_{i}\sim\mathcal{N}(t_{i}+40,0.1)$.
All models were trained with the same hyperparameters as in Table˜12. The predicted intensity functions for a single sequence can be seen in Fig.˜11, and the likelihood ratio between the trained models and the ground truth process on held-out test sequences can be found in Table˜13. We can see that both S2P2 and AttNHP do very well, both qualitatively and quantitatively. This is expected as both architectures are well suited for long-range dependencies while still being continuous-time models, allowing for expressive intensity functions.
Table 13: Likelihood ratios between models and ground truth process for held-out data on long-range experiment.
Model Ground Truth RMTPP SAHP THP IFTPP MHP NHP AttNHP S2P2
Lik. Ratio 100% 80.4% 81.0% 94.0% 88.0% 87.3% 87.9% 99.7% 97.8%
Figure 11: Synthetic long-range experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence. The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
D.6 Ablation for Different S2P2 Variants
@@
We perform an ablation study of different model variants that we proposed on all datasets and summarize the results in Table˜14. We train EHRSHOT using 10% of its training data because larger dataset scale requires more training time (but use the original validation and test sets for model selection and reporting results). Forward and backward discretization are very close in performance, with backwards discretization having a slight edge. Models that are input-dependent achieve better performance on most datasets, although on certain datasets input dependence appears to harm performance. It is an interesting direction for future work to explore theoretically and empirically when each of these variants is best. We select backward discretization with input dependence for the results in the main paper.
@@
Table 14: Ablation for different model variants log-likelihood ($L$). ID stands for input-dependent, see Section˜$3.$ Backward and Forward respectively refer to using $\mathbf{u}_{t_{i-1}}$ and $\mathbf{u}_{t_{i}-}$ (i.e., the previous right limit or current left limit), see Section˜B.4.
Dataset Model variant $L$ Arrival time LL Mark LL conditioned on time
@@
D.7 Model Calibration
To further probe the models, we evaluate the calibration of MTPPs, as proposed by Bosser and Taieb [2023]. Calibration has a different focus than log-likelihood-based or accuracy-based evaluation. Calibration instead describes how well the uncertainty in the model is reflective of the observed data. It is important to note, however, a model can achieve perfect calibration simply by predicting the marginal distribution. Better calibration therefore does not necessarily indicate better predictive performance — only better calibrated errors — and so should be taken in context with the performance under other metrics. We provide summarized statistics for both probabilistic calibration error (PCE) for time calibration and expected calibration error (ECE) for mark calibration in Table˜15, and visualize the calibration curves in Figs.˜12 and 13.
We see that, on the whole, MTPP models produce fairly well-calibrated predictions. IFTPP is the best calibrated of the models, this may be as a result of having parametric distributions for inter-arrival time (although IFTPP does fail on some datasets such as Retweet). The S2P2 is particularly well calibrated in time (PCE) among intensity-based methods, suggesting again that our S2P2 is capturing time dependencies better than other models. It is also the second-best calibrated on mark prediction (ECE) on average.